Cynthia Besch wants to buy a rug for a room that is 18 ft wide and 31 ft long. She wants to leave a uniform strip of floor around the rug. She can afford to buy 300 square feet of carpeting. What dimensions should the rug​ have? The dimensions of the rug should be blank ft.
Found 2 solutions by ikleyn, jorel1380:
Answer by ikleyn(52778)

.
It was just solved in
https://www.algebra.com/algebra/homework/quadratic/Quadratic_Equations.faq.question.1066615.html

https://www.algebra.com/algebra/homework/quadratic/Quadratic_Equations.faq.question.1066615.html



Answer by jorel1380(3719)

The room is 18 x 31 feet, yielding an area of 558 sq.ft. Let w be the width of the strip of bare floor around the carpet. Then:
(31-2w)(18-2w)=300
558-98w+4w²=300
4w²-98w+258=0
2w²-49w+129=0
(2w-43)(w-3)=0
w=43/2 or 3
Picking the most logical choice, we choose 3 ft. as the width of the strip, giving us a rug of 31-6 x 18-6, or 25' x 12' ft (300 sq.ft.). ☺☺☺☺